The length of a rectangular garden is 4m more than its width. Determine the dimensions of the garden if the area is 117m^
Answer by DAVKAY31(2)

l = 4+w
w= w
(4=w)(w) =117
4w +w^2 =117

w^2 +4w -117=0
(w + 13)(w-9) =0

w + 13 = 0 w - 9 = 0
w =-13 w = 9

w = 9
l = 13 (4 + 9)